### v1

@@ 5. Positive dimension
$\forall y\in X,\quad\mu^{*n}*\delta_{x}(B_{\rho}y)\ll\rho^{\kappa}.$
The case $d=1$ corresponds to [7, Proposition 3.1]. For arbitrary $d$, the argument of [7] goes through with a few adaptations to deal with the rotation component $K^{\prime}$. We provide the proof for completeness.
@@
(5.1) $\mu^{*n}*\delta_{x}(B_{\rho}y)>\rho^{\kappa}.$
Let $\alpha=\frac{1}{10(\ell+1)}$ and $m=\lfloor\alpha\lvert\log\rho\rvert\rfloor$. Write
@@
(5.2) $\mu^{*m}\{g:gz\in B_{\rho}y\}\geq\rho^{2\kappa}.$
On the other hand, by the large deviation principle, there exists $\varepsilon=\varepsilon(\mu)>0$ such that for $\rho\lll 1$,
(5.3) $\mu^{*m}\{g\,:\,\log\mathtt{r}_{g}\in[-(\ell+1)m,-(\ell-1)m]\}\geq 1-\rho^{\alpha\varepsilon}.$
Furthermore, considering $\gamma=\gamma(\mu)>0$ as in Lemma˜3.5, we have for all $\rho\lll_{\kappa}1$,
(5.4) $\mu^{*m}\{g\,:\,\lVert\mathtt{b}_{g}\rVert\leq\rho^{-4\gamma^{-1}\kappa}\}\geq 1-\rho^{3\kappa}.$
@@
(5.5) $\mu^{*m}(E)\geq\frac{\rho^{2\kappa}-\rho^{\alpha\varepsilon}-\rho^{3\kappa}}{|I|}\geq\rho^{10d^{2}C\kappa},$
@@
(5.6) $\operatorname{dist}(z,g_{1}^{-1}g_{2}z)\ll\lVert\operatorname{Ad}(g_{1}^{-1})\rVert\rho\ll\rho^{1/2}.$
@@
$g_{1}^{-1}g_{2}=u(-\mathtt{b}_{g_{2}})hu(\mathtt{b}_{g_{2}})\quad\text{where }h=u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}_{g_{2}}^{-1})k_{g_{1}}k_{g_{2}}^{-1}.$
To deduce an estimate on the injectivity radius of $X$ at $z$, we now show that $g_{1},g_{2}$ can be chosen so that $g_{1}^{-1}g_{2}$ is not too close to $\operatorname{Id}$, but still at a distance less than a small power of $\rho$ from one another. First, using the non-concentration estimate Lemma˜3.5 (ii) and (5.5), we can choose $g_{1},g_{2}\in E$ such that
(5.7) $\lVert\mathtt{b}_{g_{1}}-\mathtt{b}_{g_{2}}\rVert\geq\rho^{11\gamma^{-1}d^{2}C\kappa},$
provided $\kappa\ll_{C}1$ to ensure that $\rho^{11\gamma^{-1}d^{2}C\kappa}\geq e^{-m}\simeq\rho^{\alpha}$ as required in Lemma˜3.5, and $\rho\lll_{\kappa}1$. For $g_{1},g_{2}\in E$ satisfying (5.7), we have
@@
(5.8) $\rho^{1/4}\ll\rho^{(11d^{2}C+4(d+1))\gamma^{-1}\kappa}\ll\operatorname{dist}(g_{1}^{-1}g_{2},\operatorname{Id})\ll\rho^{(C-4(d+1)\gamma^{-1})\kappa},$
@@
By the effective recurrence statement from Proposition˜4.1, this is absurd if $n-m\ggg{\lvert\log\operatorname{inj}(x)\rvert}$. The proof of the proposition is complete. ∎
@@ 6.1. Non-concentration inequalities
The strategy to prove Proposition˜6.2 is to show that convolution by $\mu$ (or a suitable power $\mu^{*n}$) improves the dimensional properties of any given Frostman measure $\nu$ on $X$. Iterating this phenomenon allows to reach high dimension. The dimensional increment property for random walks is rooted in the following key observation:
$\mu^{*n}*\nu(B_{\rho}x)=\int_{G}\nu(g^{-1}B_{\rho}x)\,\,\mathrm{d}\mu^{*n}(g)$
and $g^{-1}B_{\rho}x$ can be seen, in some exponential chart, as a Euclidean box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$, varying randomly with $g\sim\mu^{*n}$. Provided this random box satisfies suitable non-concentration properties, we can then derive a small dimensional increment via a multislicing theorem (which itself boils down to the sum product phenomenon).
The required non-concentration concerns the partial flag carrying the box. Let us see what it is in our setting. Consider the weight spaces decomposition $\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{+}$ for $A^{\prime}$. More precisely, $\mathfrak{g}_{+},\mathfrak{g}_{-}$ are respectively the Lie algebras of $U$ and $U^{-}$, where $U^{-}$ denotes the transpose of $U$, and $\mathfrak{g}_{0}$ is their orthogonal complement in $\mathfrak{g}$. Recalling $g=k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})$ and the norm on $\mathfrak{g}$ is $\operatorname{Ad}(K^{\prime})$-invariant, the box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$ can be written
@@
(6.3) $\overset{L_{g}}{\simeq}B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}}_{\mathtt{r}^{-1}_{g}\rho}+B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}}_{\rho}+B^{\mathfrak{g}}_{\mathtt{r}_{g}\rho}$
where $\mathfrak{g}_{\leq 0}:=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}$, $L_{g}=O((1+\|\mathtt{b}_{g}\|)^{d+1})$, and the notation $A\overset{L}{\simeq}B$ means that $A$ can be covered by less than $L$ additive translates of $B$, and conversely. Note the norm $\|\mathtt{b}_{g}\|$ is controlled via Lemma˜3.5 (i). We are left to examine the non-concentration properties of the partial flag given by $V_{1}(g):=\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}$ and $V_{2}(g):=\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}$ as $g$ varies with law $\mu^{*n}$.
In [7] about the case $d=1$ (as well as in [6]), a similar approach is exploited, but the non-concentration at disposal therein is very strong, namely: for $i=1,2$, any subspace $W\subseteq\mathfrak{g}$ with $\dim V_{i}+\dim W=\dim\mathfrak{g}$, for $\mu^{*n}$-many $g$, we have $V_{i}(g)\cap W=\{0\}$ with a large angle between $V_{i}(g)$ and $W$. This non-concentration requirement is natural, as it is the hypothesis of the projection theorems à la Bourgain which are at the heart of the multislicing estimates from [6]. Unfortunately, such property fails for $d\neq 1$, as we see in the next lemma.
@@
Assume $d\geq 2$. Let $W=\{\,M\in\mathfrak{g}\,:\,Me_{1}=0\,\}$ be the subspace of matrices in $\operatorname{M}_{d}(\mathbb{R})$ with zero trace and null first column. Then $\operatorname{codim}W=d+1>\dim\mathfrak{g}_{-}$ but for every $g\in G$, we have
@@
$\mathfrak{g}_{-}=\oplus_{j=1}^{d}\mathbb{R}E_{d+1,j}=\begin{pmatrix}0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ *&*&\cdots\,*&0\end{pmatrix}$.
Therefore, $\operatorname{Ad}(g)\mathfrak{g}_{-}$ corresponds to the collection of endomorphisms of $\mathbb{R}^{d+1}$ which are $2$-step nilpotent and with image in $g\mathbb{R}e_{d+1}$. Consider $m,m^{\prime}\in\operatorname{Ad}(g)\mathfrak{g}_{-}$ non colinear. As $me_{1},m^{\prime}e_{1}$ are colinear, there must exist $(t,t^{\prime})\in\mathbb{R}^{2}\smallsetminus\{0\}$ with $(tm+t^{\prime}m^{\prime})e_{1}=0$, whence $tm+t^{\prime}m^{\prime}\in W$. This justifies that any $2$-dimensional subspace of $\operatorname{Ad}(g)\mathfrak{g}_{-}$ intersects $W$, whence $\dim(\operatorname{Ad}(g)\mathfrak{g}_{-}\cap W)\geq d-1$. ∎
In this section, we consider an arbitrary $d\geq 1$ and show that the random subspaces $V_{1}(g),V_{2}(g)$ where $g\sim\mu^{*n}$ still satisfy a weak form of non-concentration. It is presented as Proposition˜6.4 below. We will explain afterward how this can be utilized to perform the dimensional bootstrap.
@@
$\|F_{1}\wedge\dots\wedge F_{k}\|:=\|v_{1}\wedge\dots\wedge v_{k}\|$
@@
Observing $\dim\mathfrak{g}=d(d+2)$, Proposition˜6.4 means that for most parameters $(g_{i})_{i=1}^{d+1}$ selected by $(\mu^{*n})^{\otimes d+1}$, we have $\mathfrak{g}=\oplus_{i}V_{1}(g_{i})\bigoplus W$, and each subspace makes a rather large angle with the complementary sum.
@@
As the map $\mathfrak{g}\mapsto\mathfrak{g},A\mapsto A^{T}$ is an isometry, the claim follows from Proposition˜6.4. ∎
We now focus on establishing Proposition˜6.4. We first reduce to a purely geometric version of that result.
@@
Proposition˜6.6 is geometric in the sense that no random variable is involved. It turns out to be equivalent to Proposition˜6.4.
Proof that Proposition˜6.4 $\iff$ Proposition˜$6.$
The direct implication is clear, therefore we assume Proposition˜6.6 and check Proposition˜6.4. We write $P(\boldsymbol{s}):=\operatorname{Ad}(u(-\boldsymbol{s}))$ for conciseness. Observe that the angle function $(\mathbb{R}^{d})^{d}\rightarrow\bigwedge^{\dim\mathfrak{g}}\mathfrak{g}\simeq\mathbb{R},(\boldsymbol{s}_{i})_{i}\mapsto\lVert P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\rVert$ is Lipschitz continuous. In view of Lemma˜2.1, it suffices to show the existence of constants $C,c>0$ depending only on $\sigma$ such that for every $r>0$,
@@
(6.4) $\|P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\|=\frac{\|P(\boldsymbol{s}_{1})v_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w\|}{\|P(\boldsymbol{s}_{1})v_{-}\|\dotsm\|P(\boldsymbol{s}_{d+1})v_{-}\|}.$
@@
(6.5) $\sigma^{\otimes d+1}\{\|P(\boldsymbol{s}_{1})v_{-}\|\dotsm\|P(\boldsymbol{s}_{d+1})v_{-}\|\geq r^{-1/2}\}\ll_{\sigma}r^{\gamma}.$
On the other hand, Proposition˜6.6 guarantees that the polynomial map $(\boldsymbol{s}_{i})_{i}\mapsto P(\boldsymbol{s}_{1})v_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w$ is non-zero. As it depends continuously on $w$ and $\operatorname{Gr}(\mathfrak{g},d)$ is compact, it must have the supremum norm on the coefficients bounded below by a constant $c_{d}>0$ depending only on $d$. Combined with Corollary˜3.4, this yields
(6.6) $\sigma^{\otimes d+1}\{\|P(\boldsymbol{s}_{1})v_{-}\wedge\dotsm\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w\|<r^{1/2}\}\ll_{\sigma}r^{\gamma}$
up to taking $\gamma$ smaller. Proposition˜6.4 follows from the combination of (6.4), (6.5), (6.6). ∎
@@
Let us check that that Proposition˜6.6 and Proposition˜6.7 are equivalent.
Proof that Proposition˜6.6 $\iff$ Proposition˜6.7.
The direct implication is clear. We establish the converse. Assume by contradiction that Proposition˜6.6 fails for some $W\in\operatorname{Gr}(\mathfrak{g},d)$. Write $Z_{G}(A^{\prime})$ the centralizer of $A^{\prime}$ in $G$. Noting that $\mathfrak{g}_{-}$ is $Z_{G}(A^{\prime})U^{-}$-invariant, we obtain for every $(g_{i})_{i}\in(UZ_{G}(A^{\prime})U^{-})^{d+1}$ that
(6.7) $\operatorname{Ad}(g_{1})\mathfrak{g}_{-}+\operatorname{Ad}(g_{2})\mathfrak{g}_{-}+\dots+\operatorname{Ad}(g_{d+1})\mathfrak{g}_{-}+W\neq\mathfrak{g}.$
@@
is preserved under the action of $\operatorname{Ad}(G)$, whence of $\operatorname{Ad}(B)$. It is moreover Zariski-closed. More precisely, it is the set of $\mathbb{R}$-points of a complete $\mathbb{R}$-variety. On the other hand, $B$ is the set of $\mathbb{R}$-points of a $\mathbb{R}$-split connected solvable linear algebraic group which acts $\mathbb{R}$-morphically on this variety. Thus, by a version of the Borel fixed point theorem ([11, Proposition 15.2]), the above set contains a fixed point for $\operatorname{Ad}(B)$. This contradicts Proposition˜6.7, thus finishing the proof of the converse implication. ∎
The advantage of reducing to Proposition˜6.7 is that it constrains $W$ to belong to a finite explicit family of subspaces.
@@
$W=\oplus_{(i,j)\in S}\mathbb{R}E_{i,j}$
where $S\subseteq\{(i,j)\,:\,1\leq i<j\leq d+1\}$ and $S$ is stable by the operations $(i,j)\mapsto(i-1,j)$ and $(i,j)\mapsto(i,j+1)$ (provided $i\geq 2$ and $j\leq d$ respectively).

### v2

@@ 6. Positive dimension
$\forall y\in X,\qquad\mu^{*n}*\delta_{x}(B_{\rho}y)\ll\rho^{\kappa}.$
The case $d=1$ corresponds to [3, Proposition 3.1]. For arbitrary $d$, the argument of [3] goes through with a few adaptations to deal with the rotation component $K^{\prime}$. We provide the proof for completeness.
@@
$\mu^{*n}*\delta_{x}(B_{\rho}y)>\rho^{\kappa}.$ (6.1)
Let $\alpha=\frac{1}{10(\ell+1)}$ where $\ell>0$ was defined in (2.1) as the top Lyapunov exponent of $\operatorname{Ad}_{\star}\mu$. Set $m=\lfloor\alpha\lvert\log\rho\rvert\rfloor$. Write
@@
$\mu^{*m}\{\,g:gz\in B_{\rho}y\,\}\geq\rho^{2\kappa}.$ (6.2)
Our first step is to see that among those $g$ taking $z$ into $B_{\rho}y$, we may impose further properties without ruining the measure estimate (precise statement in (6.5) below). First, by the large deviation principle, there exists $\varepsilon=\varepsilon(\mu)>0$ such that for $\rho\lll 1$,
$\mu^{*m}\{\,g\,:\,\log\mathtt{r}_{g}\in[-(\ell+1)m,-(\ell-1)m]\,\}\geq 1-\rho^{\alpha\varepsilon}.$ (6.3)
Furthermore, considering $\gamma=\gamma(\mu)>0$ as in Lemma˜3.3, we have for all $\rho\lll_{\kappa}1$,
$\mu^{*m}\{\,g\,:\,\lVert\mathtt{b}_{g}\rVert\leq\rho^{-4\gamma^{-1}\kappa}\,\}\geq 1-\rho^{3\kappa}.$ (6.4)
@@
$\mu^{*m}(E)\geq\frac{\rho^{2\kappa}-\rho^{\alpha\varepsilon}-\rho^{3\kappa}}{|I|}\geq\rho^{10d^{2}C\kappa},$ (6.5)
@@
$\operatorname{dist}(z,g_{1}^{-1}g_{2}z)\ll\lVert\operatorname{Ad}(g_{1}^{-1})\rVert\rho\ll\rho^{1/2}.$ (6.6)
@@
$g_{1}^{-1}g_{2}=u(-\mathtt{b}_{g_{2}})hu(\mathtt{b}_{g_{2}})\qquad\text{where }h=u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}_{g_{2}}^{-1})k_{g_{1}}k_{g_{2}}^{-1}.$
To deduce from (6.6) an estimate on the injectivity radius of $X$ at $z$, we now show that $g_{1},g_{2}$ can be chosen so that the distance between $g_{1}^{-1}g_{2}$ and $\operatorname{Id}$ is much bigger than $\rho^{1/2}$ (say at least $\rho^{1/4}$), but still smaller than a power of $\rho$. Assume $\gamma\lll 1$ so that the non-concentration estimate from Lemma˜4.6 (i) holds. Combining it with (6.5), we can choose $g_{1},g_{2}\in E$ such that
$\lVert\mathtt{b}_{g_{1}}-\mathtt{b}_{g_{2}}\rVert\geq\rho^{11\gamma^{-1}d^{2}C\kappa},$ (6.7)
provided $\kappa\ll_{C}1$ to ensure that $\rho^{11\gamma^{-1}d^{2}C\kappa}\geq e^{-m}\simeq\rho^{\alpha}$ as required in Lemma˜4.6, and $\rho\lll_{\kappa}1$. For $g_{1},g_{2}\in E$ satisfying (6.7), we have
@@
$\rho^{1/4}\ll\rho^{(11d^{2}C+4(d+1))\gamma^{-1}\kappa}\ll\operatorname{dist}(g_{1}^{-1}g_{2},\operatorname{Id})\ll\rho^{(C-4(d+1)\gamma^{-1})\kappa},$ (6.8)
@@
By the effective recurrence statement from Proposition˜5.1, this is absurd if $n-m\ggg{\lvert\log\operatorname{inj}(x)\rvert}$. The proof of the proposition is complete. ∎
@@ 7. Dimensional Bootstrap
Remark.
The proof of Proposition˜7.2 yields a slightly more precise lower bound requirement on $n$: we only need $n\geq M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$ where $M\ggg_{\kappa}1$ and $A\ggg 1$. See the end of §7.4.
The strategy to prove Proposition˜7.2 is to show that convolution by $\mu$ (or a suitable power $\mu^{*n}$) improves the dimensional properties of any given Frostman measure $\nu$ on $X$. Iterating this phenomenon allows to reach high dimension. The dimensional increment property for random walks is rooted in the following key observation:
$\mu^{*n}*\nu(B_{\rho}x)=\int_{G}\nu(g^{-1}B_{\rho}x)\,\,\mathrm{d}\mu^{*n}(g)$
and $g^{-1}B_{\rho}x$ can be seen, in some exponential chart, as a Euclidean box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$ varying randomly with $g\sim\mu^{*n}$. Provided this random box satisfies suitable non-concentration properties, we can then derive a small dimensional increment via a multislicing theorem (which itself boils down to the sum product phenomenon). In §7.1, we study the non-concentration properties at our disposal. In §7.2, we develop the relevant multislicing machinery in an abstract Euclidean setting. In §7.3, we present suitable linearizing charts to allow application in the context of our homogeneous space $X$. In §7.4 we put all these results together to obtain Proposition˜7.2.
@@ 7.1. Non-concentration inequalities
We establish non-concentration properties for the partial flag carrying the box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$ as $g$ varies randomly with law $\mu^{*n}$.
First, let us see what this partial flag is explicitly. Consider the weight spaces decomposition $\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{+}$ with respect to $A^{\prime}$. More precisely, $\mathfrak{g}_{+},\mathfrak{g}_{-}$ are respectively the Lie algebras of $U$ and $U^{-}$, where $U^{-}$ denotes the transpose of $U$, and $\mathfrak{g}_{0}$ is their orthogonal complement in $\mathfrak{g}$. Recalling $g=k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})$ and the norm on $\mathfrak{g}$ is $\operatorname{Ad}(K^{\prime})$-invariant, and assuming $\mathtt{r}_{g}<1$, the box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$ can be written
$\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$ $=\operatorname{Ad}(u(-\mathtt{b}_{g}))\operatorname{Ad}(a(\mathtt{r}_{g}))B^{\mathfrak{g}}_{\rho}$
@@
[…]
$\overset{L_{g}}{\simeq}B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}}_{\mathtt{r}^{-1}_{g}\rho}+B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}}_{\rho}+B^{\mathfrak{g}}_{\mathtt{r}_{g}\rho}$ (7.3)
where $\mathfrak{g}_{\leq 0}:=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}$, $L_{g}=O(\|\operatorname{Ad}(u(\pm\mathtt{b}_{g}))\|)^{\dim\mathfrak{g}}=(2+\|\mathtt{b}_{g}\|)^{O(1)}$, and the notation $A\overset{L}{\simeq}B$ means that $A$ can be covered by less than $L$ additive translates of $B$, and conversely. Note the norm $\|\mathtt{b}_{g}\|$ is controlled via Lemma˜3.3. We are left to examine the non-concentration properties of the partial flag given by
$V_{1}(g):=\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}\qquad\text{ and }\qquad V_{2}(g):=\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}$
as $g$ varies with law $\mu^{*n}$.
In [3] about the case $d=1$ (as well as in [4]), a similar approach is exploited, but the non-concentration at disposal therein is very strong, namely: for $i=1,2$, any subspace $W\subseteq\mathfrak{g}$ with $\dim V_{i}+\dim W=\dim\mathfrak{g}$, for $\mu^{*n}$-many $g$, we have $V_{i}(g)\cap W=\{0\}$ with a large angle between $V_{i}(g)$ and $W$. This non-concentration requirement appeared naturally, as it is the hypothesis of the projection theorems [13, 29] that were previously put into the multislicing machinery of [4]. Unfortunately, such property fails for $d\neq 1$, as we see in the next lemma.
@@
Assume $d\geq 2$. Let $W=\{\,M\in\mathfrak{g}\,:\,Me_{1}=0\,\}$ be the subspace of matrices in $\operatorname{M}_{d+1}(\mathbb{R})$ with zero trace and null first column. Then $\operatorname{codim}_{\mathfrak{g}}W=d+1>\dim\mathfrak{g}_{-}$ but for every $g\in G$, we have
@@
Here and below, we denote by $(e_{1},\dotsc,e_{d+1})$ the standard basis of $\mathbb{R}^{d+1}$.
@@
$\mathfrak{g}_{-}=\operatorname{span}\{\,E_{d+1,j}\,:\,1\leq j\leq d\,\}=\begin{pmatrix}0&\cdots&0&0\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&0\\ *&\cdots&*&0\end{pmatrix}$.
Therefore, for any $g\in G$, $\operatorname{Ad}(g)\mathfrak{g}_{-}$ corresponds to the collection of endomorphisms of $\mathbb{R}^{d+1}$ which are $2$-step nilpotent and with image in $g\mathbb{R}e_{d+1}$. Consider $m,m^{\prime}\in\operatorname{Ad}(g)\mathfrak{g}_{-}$ non colinear. As $me_{1},m^{\prime}e_{1}$ are colinear, there must exist $(t,t^{\prime})\in\mathbb{R}^{2}\smallsetminus\{0\}$ with $(tm+t^{\prime}m^{\prime})e_{1}=0$, whence $tm+t^{\prime}m^{\prime}\in W$. This justifies that any $2$-dimensional subspace of $\operatorname{Ad}(g)\mathfrak{g}_{-}$ intersects $W$, whence $\dim(\operatorname{Ad}(g)\mathfrak{g}_{-}\cap W)\geq d-1$. ∎
In this section, we consider an arbitrary $d\geq 1$ and show that the random subspaces $V_{1}(g),V_{2}(g)$ where $g\sim\mu^{*n}$ still satisfy a weak form of non-concentration. It is presented as Proposition˜7.4 below. We will explain afterwards how this can be utilized to perform the dimensional bootstrap.
@@
$\lVert F_{1}\wedge\dots\wedge F_{k}\rVert:=\lVert v_{1}\wedge\dots\wedge v_{k}\rVert$
@@
Observing $\dim\mathfrak{g}=d(d+2)$, Proposition˜7.4 means that for most parameters $(g_{i})_{i=1}^{d+1}$ selected by $(\mu^{*n})^{\otimes d+1}$, we have $\mathfrak{g}=\bigoplus_{i}V_{1}(g_{i})\oplus W$, and each subspace makes a rather large angle with the complementary sum.
@@
As the map $\mathfrak{g}\mapsto\mathfrak{g},A\mapsto A^{T}$ is an isometry, the claim follows from Proposition˜7.4. ∎
We now focus on establishing Proposition˜7.4. We first reduce to a purely geometric version of that result.
@@
Proposition˜7.6 is geometric in the sense that no random variable is involved. It turns out to be equivalent to Proposition˜7.4.
Proof that Proposition˜7.4 $\iff$ Proposition˜7.6.
The direct implication being clear, we assume Proposition˜7.6 and check Proposition˜7.4. We write $P(\boldsymbol{s}):=\operatorname{Ad}(u(-\boldsymbol{s}))$ for conciseness. Observe that the angle function $(\mathbb{R}^{d})^{d}\rightarrow\bigwedge^{\dim\mathfrak{g}}\mathfrak{g}\simeq\mathbb{R},(\boldsymbol{s}_{i})_{i}\mapsto\lVert P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\rVert$ is Lipschitz continuous. In view of Lemma˜3.2, it suffices to show the existence of constants $C,c>0$ depending only on $\sigma$ such that for every $r>0$,
@@
$\|P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\|=\frac{\|P(\boldsymbol{s}_{1})v_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w\|}{\|P(\boldsymbol{s}_{1})v_{-}\|\dotsm\|P(\boldsymbol{s}_{d+1})v_{-}\|}.$ (7.4)
@@
$\sigma^{\otimes d+1}\left\{\|P(\boldsymbol{s}_{1})v_{-}\|\dotsm\|P(\boldsymbol{s}_{d+1})v_{-}\|\geq r^{-1/2}\right\}\ll_{\sigma}r^{\gamma}.$ (7.5)
On the other hand, Proposition˜7.6 guarantees that the polynomial map $(\boldsymbol{s}_{i})_{i}\mapsto P(\boldsymbol{s}_{1})v_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w$ is non-zero. As it depends continuously on $w$ and $\operatorname{Gr}(\mathfrak{g},d)$ is compact, it must have the supremum norm on the coefficients bounded below by a constant $c_{d}>0$ depending only on $d$. Combined with Corollary˜4.5, this yields
$\sigma^{\otimes d+1}\left\{\|P(\boldsymbol{s}_{1})v_{-}\wedge\dotsm\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w\|<r^{1/2}\right\}\ll_{\sigma}r^{\gamma}$ (7.6)
up to taking $\gamma$ smaller. Proposition˜7.4 follows from the combination of (7.4), (7.5), (7.6). ∎
@@
Let us check that Proposition˜7.6 and Proposition˜7.7 are equivalent.
Proof that Proposition˜7.6 $\iff$ Proposition˜$7.$
The direct implication is clear. We establish the converse. Assume by contradiction that Proposition˜7.6 fails for some $W\in\operatorname{Gr}(\mathfrak{g},d)$. Write $Z_{G}(A^{\prime})$ the centralizer of $A^{\prime}$ in $G$. Noting that $\mathfrak{g}_{-}$ is $Z_{G}(A^{\prime})U^{-}$-invariant, we obtain for every $(g_{i})_{i}\in(UZ_{G}(A^{\prime})U^{-})^{d+1}$ that
$\operatorname{Ad}(g_{1})\mathfrak{g}_{-}+\operatorname{Ad}(g_{2})\mathfrak{g}_{-}+\dots+\operatorname{Ad}(g_{d+1})\mathfrak{g}_{-}+W\neq\mathfrak{g}.$ (7.7)
@@
is preserved under the action of $\operatorname{Ad}(G)$, whence of $\operatorname{Ad}(B)$. It is moreover Zariski-closed. More precisely, it is the set of $\mathbb{R}$-points of a complete $\mathbb{R}$-variety. On the other hand, $B$ is the set of $\mathbb{R}$-points of a $\mathbb{R}$-split connected solvable linear algebraic group which acts $\mathbb{R}$-morphically on this variety. Thus, by a version of the Borel fixed point theorem ([11, Proposition 15.2]), the above set contains a fixed point for $\operatorname{Ad}(B)$. This contradicts Proposition˜7.7, thus finishing the proof of the converse implication. ∎
The advantage of reducing to Proposition˜7.7 is that it constrains $W$ to belong to a finite explicit family of subspaces.
@@
$W=\operatorname{span}\{\,E_{i,j}\,:\,(i,j)\in S\,\}$
for some subset $S\subseteq\{\,(i,j)\,:\,1\leq i<j\leq d+1\,\}$ that is stable under the operations $(i,j)\mapsto(i-1,j)$ and $(i,j)\mapsto(i,j+1)$ (provided that $i\geq 2$ and $j\leq d$ respectively).